X+10Y+Z=52
5X+Y+4Z=15
X+2y-3Z=12

Answer by Edwin McCravy(20086)

X+10Y+Z=52
5X+Y+4Z=15
X+2y-3Z=12


There are many differen ways to solve that syatem and since you
didn't specify which way you are studying, I'll just give the
solution.  If you post again telling which method you are
studying we will show you step by step how to do it.  

Solution:  (X,Y,Z) = (2,5,0) 

Edwin